Problem 7-12

You are 25 years old and inherit $60,000 from your grandmother. If you wish to purchase a $107,450 boat to celebrate your 35th birthday, what compound annual rate of return must you earn? Use Appendix A to answer the question. Round your answer to the nearest whole number.
Answer #1

Ans 6.00%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
107450 / 60000 = (1 + r%)^10
1.79083333333= (1 + r%)^10
r = 6.00%